OpenStudy (anonymous):

The sum of two numbers is 70 and the difference is 11. What are the numbers

OpenStudy (anonymous):

\[x+y=70\]\[x-y=11\] If you add the two system of equations, you get\[2x = 81\]With that, you can solve for what x is\[x=\frac{81}{2}\] You can now plug that x into either equations and solve for y. Let's use the first one.\[\frac{81}{2}+y=70\]Now we have to solve for y. Subtract 81/2 from both sides.\[y = \frac{59}{2}\] Now we have both x and y\[x=\frac{81}{2}, y=\frac{59}{2}\] You can plug both those numbers into your 2nd equation and verify that \[x-y=11\]I'll leave you to do that.